if every fourth trick-or-treater is a vampire, and every 6th, robot, how many of each will you see if seventy-five trick -or treaters come to your door?
Answer by stanbon(75887)

if every fourth trick-or-treater is a vampire, and every 6th, robot, how many of each will you see if seventy-five trick -or treaters come to your door?
---
# of vampires: 75/4 = 18 + 3
Ans: 18
---
# of robots: 75/6 = 12 + 3
Ans: 12
============================
Cheers,
Stan H.